# ReDit: Reward Dithering for Improved LLM Policy Optimization

### v2

@@ 3.2 Theoretical Principles to Address the Limitations of Discrete Rewards
Figure 5: Test accuracy across datasets on Qwen2.5-7B-Instruct and Qwen2.5-VL-7B-Instruct. The horizontal dashed line marks ReDit’s performance at 1000 steps, which GRPO fails to match even after 9000 steps.
4 Reward Dithering (ReDit)
The crucial modification introduced by ReDit lies in how the advantage term $\hat{A}_{i,t}^{\text{GRPO}}$ within Eq. (2) is computed. Instead of directly using the raw discrete rewards $r_{i}=r(o_{i})$ obtained for each sampled output $o_{i}$ in the batch $\{o_{i}\}_{i=1}^{G}$ (line 3 in Algorithm 1), we first compute smoothed rewards $\tilde{r}_{i}$. This is done by adding independently sampled zero-mean perturbation $\epsilon_{i}$ (e.g., from $\mathcal{N}(0,\sigma^{2})$ or $\mathcal{U}[-a,a]$) to each discrete reward (line 6 in Algorithm 1):
$\tilde{r}_{i}=r_{i}+\epsilon_{i}$ (5)
$\hat{A}_{i,t}^{\text{GRPO}}$ $\propto\frac{r_{i}-\text{mean}(\{{r}_{k}\}_{k=1}^{G})}{\text{std}(\{{r}_{k}\}_{k=1}^{G})}$ (6)
[…]
Thus, the relative standing of each output $o_{i}$ within the batch, which informs its advantage $\hat{A}_{i,t}^{\text{GRPO}}$ used in Eq. (2), is determined by the continuous smoothed reward $\tilde{r}_{i}$ rather than the discrete $r_{i}$. This substitution transforms the optimization landscape. By introducing continuous variations via $\tilde{r}_{i}$, the added noise provides informative, non-zero gradients even when discrete rewards $r_{i}$ are sparse or identical within a batch, mitigating gradient vanishing. It also dampens the sharp changes in expected advantage resulting from small policy shifts affecting discrete outcomes, thus reducing the likelihood of gradient explosion. This overall smoothing effect facilitates a more stable gradient flow, enabling more robust and efficient optimization of the policy $\pi_{\theta}$ using the GRPO objective (line 8 in Algorithm 1).
@@ 5.1 Datasets and Setup
Datasets. Our dataset selection and setup largely follow the methodology of (Shao et al., 2024), primarily to assess the mathematical reasoning capabilities of the models. This encompasses mathematical problem-solving datasets such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), as well as the multimodal geometric reasoning dataset Geometry3K (Lu et al., 2021). Each dataset provides distinct training and test splits, which we utilize accordingly for model training and subsequent evaluation. See the Appendix D.1 for details of the dataset.
@@
Initial Policy. To rigorously assess the effectiveness of ReDit without confounding factors introduced by supervised fine-tuning (SFT), we initialized our experiments directly with instruct models without any additional SFT training. Previous research by Shao et al. (2025) demonstrated that even random rewards can enhance performance for Qwen models. Therefore, we conducted comprehensive evaluations across a diverse set of instruction-tuned models, including Qwen2.5-7B-Instruct, Qwen2.5-VL-7B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Ministral-8B-Instruct-2410, and Mistral-7B-Instruct-v0.3, to establish the generalizability of ReDit.
Other Training Settings. For parameter-efficient fine-tuning, we employed Low-Rank Adaptation (LoRA) (Hu et al., 2022). Our implementation leverages the official GRPO implementation within the TRL library (von Werra et al., 2020). Specific configurations for LoRA and GRPO parameters are detailed in the Appendix D.3. Model evaluation was conducted using the OpenCompass (Contributors, 2023). All experiments were executed on single NVIDIA H20 GPU.
@@ 5.2 Main Results
In our main experiments, we validate the effectiveness of our proposed ReDit. For these experiments, we primarily use either a uniform smoothing kernel with radius $a=0.05$ or a Gaussian smoothing kernel with standard deviation $\sigma=a/\sqrt{3}$. More experimental results can be found in the Appendix E.
Accelerated Convergence Across Datasets and LLMs. We demonstrate that integrating our proposed method, ReDit, with GRPO substantially accelerates convergence and improves final performance across a wide range of datasets (Fig. 5) and LLMs, including Llama-3.2-3B, Llama-3.1-8B, Ministral-8B and Mistral-7B (Fig. 6). On all tested models, both Gaussian and uniform variants of ReDit enable GRPO to reach a competitive performance level within merely 1000 training steps. Notably, this performance already surpasses that of the baseline GRPO trained for the full 9000 steps. Consequently, ReDit not only enhances training efficiency but also leads to superior final accuracy. The Gaussian variant, in particular, consistently yields the strongest results and promotes more stable training trajectories with lower volatility compared to the baseline.
Table 1: Test accuracy comparison across datasets for original Backbone, GRPO, and ReDit represents percentage point improvement of the superior ReDit variant over the GRPO.
@@
w/ ours(Gauss) 90.76 52.55 44.67
w/ ours(Uniform) 90.46 51.96 44.36
@@
Table 2: Comparison of the accuracy for different baselines under 9000 steps on GSM8K.
Name DAPO DR.GRPO REINFORCE++
Baseline 87.52 86.13 86.25
w/ ours(Gauss) 89.34 87.69 87.96
w/ ours(Uniform) 88.57 87.34 87.59
$\Delta$ +1.82 +1.56 +1.71
Optimal Performance with Scheduled Perturbation. We further investigate convergence behavior under various scheduled perturbation schemes: SquareRoot, Cosine, and CosineReverse perturbations. These schedules dynamically adjust perturbation variance throughout training, potentially benefiting model learning. Fig. in the Appendix E.7 illustrates the different perturbation schedules, while Fig. 7 presents their performance. Compared to standard GRPO, ReDit achieves both faster convergence and superior final performance, with the CosineReverse perturbation schedule yielding particularly strong results. Additional details are provided in the Appendix E.7.
Figure 7: CosineReverse achieves the best performance.
Figure 8: Accuracy of different GRPO variants tested on the GSM8K dataset. The horizontal dashed line highlights the performance of using ReDit at about 1000 training steps, and even after 9000 steps, its accuracy is comparable to the baseline.
@@ 5.3 Ablation Studies
Figure 9: Appropriate perturbation achieves the best performance.
@@
Comparison with Direct Gradient Manipulation Baselines. We benchmark ReDit against established techniques that directly address gradient instability: Gradient Clipping (Zhang et al., 2020), which mitigates exploding gradients, and Dynamic Sampling (Yu et al., 2025), which alleviates vanishing gradients. The objective is to compare our ReDit approach with methods that operate directly on the gradient signal. As illustrated in Figure 11, ReDit substantially outperforms both baseline methods. We attribute this performance gap to the inherent limitations of these heuristics. Gradient Clipping, for instance, crudely truncates gradient magnitudes, a non-principled operation that can introduce significant estimation bias. Conversely, while Dynamic Sampling can be effective for vanishing gradients, it offers no mechanism to prevent gradients from exploding. In contrast, ReDit stabilizes the training process by smoothing the reward, which provides a more principled solution to prevent both gradient vanishing and explosion, thereby leading to more efficient and effective training.
@@ 6 Theoretical Insights
Problem Setup. Our analysis uses a simplified RL framework (from Eq. (1)) focusing on binary rewards $R(q,o)\in\{0,1\}$ for complete outputs (e.g., GRPO (Shao et al., 2024)), not token-level rewards. We investigate how Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ injection improves convergence. The perturbed objective is:
@@
Proposition 6.1 (Unbiased estimate of gradient).
@@
Remark. Proposition  6.1 provides theoretical proof that introducing Gaussian noise perturbations into the discrete reward signal preserves the unbiased nature of the policy gradient estimate. This means that, under the perturbed reward, the expected direction of the policy update is consistent with the original objective being optimized. Maintaining this unbiased nature ensures that the injected noise does not introduce systematic biases into the learning dynamics, thus providing a theoretical basis for the empirical observation that our approach helps to consistently improve performance. See Appendix  B.1 for a detailed proof.
Proposition 6.2 (Introducing the variance of gradient estimation).
@@
Remark. In Proposition 6.2, we analyze how Gaussian reward perturbations affect the variance of policy gradient estimates. Adding Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ to the reward introduces a ”gradient noise” component proportional to $\epsilon\cdot\nabla_{\theta}\log\pi_{\theta}(o|q)$ in the gradient estimate. The increased variance has significant optimization benefits: Mitigate vanishing gradients: Gradient noise provides consistent stochastic updates even when the original gradient terms are small or vanishing, thus helping to avoid flat regions. Avoid exploding gradients: The randomness induced by the noise enables the optimization trajectory to probabilistically bypass unstable regions of high curvature. Furthermore, the noise variance $\sigma$ can be adjusted to control the magnitude of the gradient noise for optimal results. This mechanism enhances the robustness of policy optimization and explains the empirical improvements observed in training stability and convergence speed from reward perturbations. For detailed derivation, see Appendix B.2.
Proposition 6.3 (Optimization time upper and lower bounds).
When we introduce Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ to the reward, it not only increases the variance of the reward, but also reduces the accuracy of the originally perfect reward function. According to (Razin et al., 2025), for the target $J(\pi_{\theta})$ (Eq. (7)), the time $t_{\gamma}$ required for the real reward $r_{G}$ to increase $\gamma$ satisfies: $\Omega\left(\left(\mathbb{E}_{q\sim p_{Q}}\left[\mathrm{var}_{o\sim\pi_{\theta}(\cdot|q)}R(q,o)\right]+\sigma^{2}\right)^{-\frac{1}{3}}\right)\leq t_{\gamma}.$ $t_{\gamma}\leq\mathcal{O}\left(\pi_{\theta(0)}(\mathbf{y}^{\gamma}|\mathbf{x})^{-1}\right).$
Remark. Proposition  6.3 analyzes the time required for a policy to reach or exceed a certain performance threshold starting from $\pi_{\theta(0)}$. Compared with optimization using the unperturbed reward, the perturbed reward introduces a larger reward variance, which can reduce the lower bound of the convergence time. Meanwhile, the perturbation introduces inaccuracies relative to the perfect reward. This inaccuracy limits the upper bound of the convergence time. This advantage arises because the random noise in the perturbation can effectively improve the perturbed rewards of non-zero probability outputs under the initial policy $\pi_{\theta(0)}$, thereby effectively encouraging broader exploration, see Fig. 1. This makes these outputs easier to perform gradient ascent, which helps to discover outputs with higher true rewards. For detailed proof, see Appendix B.3.

### v4

@@ 5 Empirical Results
Table 1: Comparison of the mean and variance of accuracy for different baselines under 9000 steps on GSM8K.
Name DAPO DR.GRPO REINFORCE++
Baseline 87.52 86.13 86.25
w/ ours(Gauss) 89.34 ($\pm$ 0.04) 87.69 ($\pm$ 0.06) 87.96 ($\pm$ 0.03)
w/ ours(Uniform) 88.57 ($\pm$ 0.01) 87.34 ($\pm$ 0.07) 87.59 ($\pm$ 0.09)
$\Delta$ +1.82 +1.56 +1.71
@@ 5.1 Datasets and Setup
Datasets. Our dataset selection and setup largely follow the methodology of shao2024deepseekmathpushinglimitsmathematical , primarily to assess the mathematical reasoning capabilities of the models. This encompasses mathematical problem-solving datasets such as GSM8K cobbe2021trainingverifierssolvemath and MATH hendrycks2021measuringmathematicalproblemsolving , as well as the multimodal geometric reasoning dataset Geometry3K lu2021intergpsinterpretablegeometryproblem . Each dataset provides distinct training and test splits, which we utilize accordingly for model training and subsequent evaluation. See the Appendix D.1 for details of the dataset.
@@
Initial Policy. To rigorously assess the effectiveness of ReDit without confounding factors introduced by supervised fine-tuning (SFT), we initialized our experiments directly with instruct models without any additional SFT training. Previous research by shao2025spuriousrewardsrethinkingtraining demonstrated that even random rewards can enhance performance for Qwen models. Therefore, we conducted comprehensive evaluations across a diverse set of instruction-tuned models, including Qwen2.5-7B-Instruct, Qwen2.5-VL-7B-Instruct qwen2025qwen25technicalreport , Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct AIAtMeta2024Llama , Ministral-8B-Instruct-2410, and Mistral-7B-Instruct-v0.3 jiang2023mistral7b , to establish the generalizability of ReDit.
Random Seeds. Our method incorporates random sampling, which can introduce variance to the experimental outcomes. To thoroughly assess the impact of this stochasticity and ensure the robustness of our results, we executed all main experiments using multiple distinct random seeds. Specifically, we selected five seeds: None (no seed), 42, 123, 888, 2025, and 9999. All metrics reported in our final results represent the mean and variance computed across these five independent runs.
Other Training Settings. For parameter-efficient fine-tuning, we employed Low-Rank Adaptation (LoRA) hu2022lora . Our implementation leverages the official GRPO implementation within the TRL library vonwerra2022trl . Specific configurations for LoRA and GRPO parameters are detailed in the Appendix D.3. Model evaluation was conducted using the OpenCompass 2023opencompass . All experiments were executed on one NVIDIA H20 GPU.
@@ 5.2 Main Results
Table 2: Test mean and variance of accuracy comparison across datasets for original Backbone, GRPO, and ReDit.
@@
w/ ours(Gauss) 90.76 ($\pm$ 0.06) 52.55 ($\pm$ 0.03) 44.67 ($\pm$ 0.03)
w/ ours(Uniform) 90.46 ($\pm$ 0.07) 51.96 ($\pm$ 0.06) 44.36 ($\pm$ 0.04)
@@
In our main experiments, we validate the effectiveness of our proposed ReDit. For these experiments, we primarily use either a uniform smoothing kernel with radius $a=0.05$ or a Gaussian smoothing kernel with standard deviation $\sigma=a/\sqrt{3}$. More experimental results can be found in the Appendix F.
Figure 5: Accuracy of different GRPO variants (DAPO, DR.GRPO, REINFORCE++) tested on the GSM8K dataset. The horizontal dashed line highlights the performance of using ReDit at about 1000 training steps, and even after 9000 steps, its accuracy is comparable to the baseline.
Accelerated Convergence Across Datasets and LLMs. We demonstrate that integrating our proposed method, ReDit, with GRPO substantially accelerates convergence and improves final performance across a wide range of datasets (Fig. 7) and LLMs, including Llama-3.2-3B, Llama-3.1-8B, Ministral-8B and Mistral-7B (Fig. 6). On all tested models, both Gaussian and uniform variants of ReDit enable GRPO to reach a competitive performance level within merely 1000 training steps. Notably, this performance already surpasses that of the baseline GRPO trained for the full 9000 steps. Consequently, ReDit not only enhances training efficiency but also leads to superior final accuracy. The Gaussian variant, in particular, consistently yields the strongest results and promotes more stable training trajectories with lower volatility compared to the baseline. We also present more experimental results in the appendix. The results on the code dataset are detailed in Appendix E.1, the results on the full parameter fine-tuning method are detailed in Appendix E.2, and the results on the Deepseek-R1 distillation model are detailed in Appendix E.3.
@@
Optimal Performance with Scheduled Perturbation. We further investigate convergence behavior under various scheduled perturbation schemes: SquareRoot, Cosine, and CosineReverse perturbations. These schedules dynamically adjust perturbation variance throughout training, potentially benefiting model learning. Fig. in the Appendix F.4 illustrates the different perturbation schedules, while Fig. 9presents their performance. Compared to standard GRPO, ReDit achieves both faster convergence and superior final performance, with the CosineReverse perturbation schedule yielding particularly strong results. Additional details are provided in the Appendix F.4.
Figure 7: Test accuracy across datasets. The horizontal dashed line marks ReDit’s performance at 1000 steps, which GRPO fails to match even after 9000 steps.
Figure 8: CosineReverse perturbation achieves the best performance.
Figure 9: Appropriate perturbation achieves the best performance.
@@ 5.3 Ablation Studies
Comparison with Direct Gradient Manipulation Baselines. We benchmark ReDit against established techniques that directly address gradient instability: Gradient Clipping Zhang , which mitigates exploding gradients, and Dynamic Sampling yu2025dapoopensourcellmreinforcement , which alleviates vanishing gradients. The objective is to compare our ReDit approach with methods that operate directly on the gradient signal. As illustrated in Figure 11, ReDit substantially outperforms both baseline methods. We attribute this performance gap to the inherent limitations of these heuristics. Gradient Clipping, for instance, crudely truncates gradient magnitudes, a non-principled operation that can introduce significant estimation bias. Conversely, while Dynamic Sampling can be effective for vanishing gradients, it offers no mechanism to prevent gradients from exploding. In contrast, ReDit stabilizes the training process by smoothing the reward, which provides a more principled solution to prevent both gradient vanishing and explosion, thereby leading to more efficient and effective training.
@@ 6 Theoretical Insights
Problem Setup. Our analysis uses a simplified RL framework (from Eq. (1)) focusing on binary rewards $R(q,o)\in\{0,1\}$ for complete outputs (e.g., GRPO shao2024deepseekmathpushinglimitsmathematical ), not token-level rewards. We investigate how Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ injection improves convergence. The perturbed objective is:
@@
Proposition 1 (Unbiased estimate of gradient).
@@
Remark. Proposition  1 provides theoretical proof that introducing Gaussian noise perturbations into the discrete reward signal preserves the unbiased nature of the policy gradient estimate. This means that, under the perturbed reward, the expected direction of the policy update is consistent with the original objective being optimized. Maintaining this unbiased nature ensures that the injected noise does not introduce systematic biases into the learning dynamics, thus providing a theoretical basis for the empirical observation that our approach helps to consistently improve performance. See Appendix  B.2 for a detailed proof.
Proposition 2 (Introducing the variance of gradient estimation).
@@
Remark. In Proposition 2, we analyze how Gaussian reward perturbations affect the variance of policy gradient estimates. Adding Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ to the reward introduces a "gradient noise" component proportional to $\epsilon\cdot\nabla_{\theta}\log\pi_{\theta}(o|q)$ in the gradient estimate. The increased variance has significant optimization benefits: Mitigate vanishing gradients: Gradient noise provides consistent stochastic updates even when the original gradient terms are small or vanishing, thus helping to avoid flat regions. Avoid exploding gradients: The randomness induced by the noise enables the optimization trajectory to probabilistically bypass unstable regions of high curvature. Furthermore, the noise variance $\sigma$ can be adjusted to control the magnitude of the gradient noise for optimal results. This mechanism enhances the robustness of policy optimization and explains the empirical improvements observed in training stability and convergence speed from reward perturbations. For detailed derivation, see Appendix B.3.